OpenStudy (sabrina1):

hi can someone plz help me solve this? i have my math exam in 2 hourS and cannot find the answer anywher ...... A sector of a circle of radius 12 miles has an area of 72 square miles. Find the central angle of this sector in degrees.

OpenStudy (anonymous):

\[\frac{x}{72}=\frac{2\pi}{144\pi}\]

OpenStudy (sabrina1):

so i just solve for x for the angle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the formula for the area of a sector is as follows: \[(\Theta \div360)\times \pi(r)^{?}\] so substituting the values would give (Theta/360) x pi(12)^2 = 72 so (Theta/360) = 72/(pi(12)^2) therefore \[\theta = (72/\pi(12)^{2})\times360\] gives you something like 57.3 degs

OpenStudy (sabrina1):

thanks really appreciate it :)